OpenStudy (anonymous):

I tried but I stuck when I simplified all the terms to 5's power of log base5 two and something~~ could someone help me ?5^2(log5 (2+X))-2=5^(X+log5(2)) solve for x.

OpenStudy (dumbcow):

\[\large 5^{2 \log_{5}(x+2)} -2 = 5^{x+\log_{5}2} \] ?

OpenStudy (anonymous):

Ya that's the question, sorry that I've no idea how to type those correctly~~

OpenStudy (dumbcow):

ok lets simplify using some exponent properties \[a^{mn} = (a^{m})^{n}\] \[a^{m+n} = a^m *a^n\] \[\rightarrow \large (5^{\log_{5}(x+2)})^{2} -2 = 5^{x}* 5^{\log_{5}2} \]

OpenStudy (dumbcow):

\[b^{\log_{b} x} = x \] when base of log matches exponential base, they cancel each other out

OpenStudy (dumbcow):

\[\rightarrow (x+2)^{2} -2 = 2*5^x\]

OpenStudy (anonymous):

Oh I get it. Thank you sooooo much!

OpenStudy (dumbcow):

your welcome...but just so you know, this cant be solved algebraically because "x" is both an exponent and part of polynomial

OpenStudy (anonymous):

mmhmm, I gonna try to work it out~~ I'm was sure but now I'm not very sure.

OpenStudy (dumbcow):

what class are you taking? you cant work it out because there is no way to isolate "x" by looking at it and checking....i see that x=0 is a solution

OpenStudy (anonymous):

I'm taking IB HL maths. It's possible to solve because the exponent suppose to be 2, it's 5(log5(2+x)) instead of the other way, sorry that was my bad, I didn't have a good look at it. Now it's so much easier. Thank you very much!!!!!!